OpenStudy (anonymous):

How do I graph this? y=-cos(@/4-pi)-1

OpenStudy (anonymous):

\[y =-\cos ( \theta \4 - \pi) -1\]

OpenStudy (anonymous):

\[y =-\cos ( \theta \4 - \pi) -1\]